(a) Find the volume of the solid generated by revolving the region bounded by the curves y = x2 and y = 4x−x2 about the line y = 6.
1
Expert's answer
2020-07-01T18:44:03-0400

"V = V1 - V2 =" "\\pi\\int^2_0 (6 - x\ufeff^2)^2 dx - \\pi\\int^2_0 (6 -(4x - x\ufeff^2))^2 dx =""= \\pi\\int^2_0 ((6 - x^2)^2 - (6 -(4x - x\ufeff^2))^2) dx = \\pi\\int^2_0 (6 - x^2 - 6 + 4x - x^2)(6 - x^2 + 6 -"

"- 4x + x^2) dx = \pi\int^2_0 (-2x^2 + 4x)(12 - 4x)dx =" "=\pi\int^2_0 (8x^3 - 16x^2 - 24x^2 + 48 x)dx = \pi (2x^4 -" "40x^3\/3 + 24x^2)|^2_0 ="

"= \pi (32 - 320/3 +96) = \pi (128 - 320/3) =" "\dfrac{(384-320)\pi}{3} = \dfrac{64\pi}{3}"